Database last updated on Thu, 29 Jan 2026 16:10:51 +0000 by localhost

Instrument constant changes for the week before 2024278 (1 sec, 0 rows)

HG burned out? (AP results when HG is ON) (0 sec, 16 rows)

Measurements and tests summary.


_jday_loc_serial
278RESOLUTE#205
(0 sec)

Dead time (0 sec)

Nothing to report.

Standard lamp R6 (0 sec)

Nothing to report.

Standard lamp R5 (1 sec)

Nothing to report.

Number of days since last dome/window cleaning (0 sec, 2 rows)

serialdays
205more than a year

Missing files (0 sec, 1 rows)

Seriallast datadays ago
20520242790-1

Run/Stop results for last week. (0 sec, 1 rows)

SerialTotalFailed 0Failed darkFailed 1Failed 2Failed 3Failed 4Failed 5Failed double
205400000000

Ozone results for good DS obs (std<2.5DU, mu<4) within 3 minutes between 2 Brewers. (1 sec)

No ozone measurements to display.

Unsorted comments (1 sec, 11470 rows)

The END (0 sec, 2 rows)

If you have questions and/or comments about this page please contact Volodya Savastiouk